zarkam21 (zarkam21):

Use the substitution method to solve the system of equations below. Enter your answer as an ordered pair. 3x-y=7 2x-2y=2

DivineSolar (divinesolar):

She is offline wait for her to get back.

OpenStudy (anonymous):

are you like stalking me now? lol @DivineSolar

DivineSolar (divinesolar):

@maddybz No, look up the definition of stalking, this is an area of my expertise.

OpenStudy (kidrah69):

\[3x-y=7 \] subtract 3x from both sides \[-y=-3x+7\] divide everything by negative y \[y=3x-7\] now use substitution and plug this in equation 2 :) \[2x-2(3x-7)=2\] solve this and get \[2x-6x+14=2\] combine like terms \[-4x=-12\] divide by -4 on both sides x=3 now plug that into 1st equation \[y=3(3)-7\] solve \[y=9-7\] y=2

zarkam21 (zarkam21):

Okay what does x = then

OpenStudy (kidrah69):

Look at the work i did its all there

zarkam21 (zarkam21):

Okay what does x = then

zarkam21 (zarkam21):

Okay what does x = then

zarkam21 (zarkam21):

Okay what does x = then

zarkam21 (zarkam21):

(3,2) ?

OpenStudy (kidrah69):

yep